An experiment is performed to determine whether the average nicotine content of one kind of cigarette exceeds that of another kind by 0.20 milligram. If =50 cigarettes of the first kind had an average nicotine content of =2.61 milligrams with a standard deviation of =0.12 milligram, whereas =40 cigarettes of the other kind had an average nicotine content of = 2.38 milligrams with a standard deviation of =0.14 milligram, test the null hypotheses - = 0.20 against the alternative hypotheses - ¹ 0.20 at the 0.05 level of significance. 


1
Expert's answer
2021-03-22T04:59:14-0400

"H_0 : μ_1 -μ_2 = 0.20 \\

H_1 : μ_1 -μ_2 > 0.20 \\

α = 0.05 \\

\bar{x_1}= 2.61 \\

s_1 = 0.12 \\

n_1 = 50 \\

\bar{x_2}= 2.38 \\

s_2 = 0.14 \\

n_2 = 40 \\

δ = 0.20"

We have independent random samples from two normal populations having mean and known variances. We want to teat the null hypothesis "\u03bc_1 -\u03bc_2 = \u03b4" , where δ is a given constant, against of the alternative "\u03bc_1 -\u03bc_2 > \u03b4" . The appropriate test statistic for this test is

Test concerning differences between means

"z = \frac{\bar{x_1}-\bar{x_2}-δ}{\sqrt{\frac{σ_1^2}{n_1} + \frac{σ_1^2}{n_2}}}"

When we deal with independent random sample from populations with unknown variances which may not even be normal, the test can still be used with "s_1" substituted for "\u03c3_1" and "s_2" substituted for "\u03c3_2" so long as both samples are large enough for the central limit theorem to be invoked.

"z = \frac{2.61-2.38-0.2}{\sqrt{\frac{(0.12)^2}{50} + \frac{(0.14)^2}{40}}} = 1.08"

Reject "H_0" if z≥1.96

Since z=1.08 < 1.96, we cannot reject the null hypothesis.